OpenStudy (anonymous):

(abc)2 = This expression without exponents is A.abcabc B.aabbcc C.a2b2c2 D.ababab

OpenStudy (anonymous):

b

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@rockstar: that would be \[a^2b^2c^2\]

OpenStudy (anonymous):

that has exponents

OpenStudy (anonymous):

Yeah, that's why your answer is wrong

OpenStudy (anonymous):

The answer should be C

OpenStudy (anonymous):

the question says without exponents dumbass

OpenStudy (anonymous):

C.a2b2c2

OpenStudy (anonymous):

I know, that's why the answer is C. a2 is the same as 2a.

OpenStudy (anonymous):

a2 =a*a 2a = a+a

OpenStudy (anonymous):

B is definitely not the answer. That has exponents.

OpenStudy (anonymous):

No, 2a = a*a

OpenStudy (anonymous):

a*a=a^2

OpenStudy (anonymous):

Rockstar, there's two people and the entire world of math against you. Your are wrong.

OpenStudy (anonymous):

(abc)^2 = (abc)(abc) = aabbbcc

OpenStudy (anonymous):

aabbcc

OpenStudy (anonymous):

(abc)2 =2a2b2c is c b is wrong

OpenStudy (anonymous):

the equation is time not power

OpenStudy (anonymous):

He didn't ask (abc)^2, he asked (abc)2

OpenStudy (anonymous):

yes rguota7 is correct

OpenStudy (anonymous):

when the number comes after the variable, it is generally assumed that the number is an exponent

OpenStudy (anonymous):

here the copy from question (abc)2 = not (abc)^2 =

OpenStudy (anonymous):

(abc)2 =2a2b2c (abc)^2=(abc)*(abc)

OpenStudy (anonymous):

nvrmind, u guys dont understand